The volume of a cube is 8 cubic centimetres. Find its total surface area
Answer by ewatrrr(24785)

 
Hi,
.
V = s^3 = 8cm^3
s^3 = 8cm^3
s = 2cm
.........
total surface area = 6s^2 = 6(2cm)^2 = 24cm^2